Should print ENTWURF over all pages, quite dark and on the bottom and with landscape orientation.

\usepackage[landscape,bottom]{draftcopy}

Test 12 This is a test page 1

Test 12 This is a test page 2

This is a testpage 1 This is a testpage 2 This is a testpage 3 This is a testpage 4